Solve the following initial value problem


Ut(x,t)=10Uxx(x,t) -10

U(-1,t)=U(1,t) Ux(-1,t)=Ux(1,t) t>0

Ux(x,0)=x+1 -1


1
Expert's answer
2021-10-20T18:33:48-0400

Solution of heat equation:

"u_n(x,t)=B_nsin(\pi nx/L)e^{kt(\pi n/L)^2}"

"k=10"

"B_n=\frac{2}{L}\int^L_0f(x)sin(\frac{\pi nx}{L})dx" , n=1,2,3,...

"f(x)=u(x,0)=10"


Then:

"-10u(-1,t)=u(1,t)" :

"10\sum B_nsin(\pi n/L)e^{10t(\pi n/L)^2}=\sum B_nsin(\pi n/L)e^{10t(\pi n/L)^2}"

"\implies sin(\pi n/L)=0 \implies L=1"


"u_x=\sum B_n \frac{\pi n}{L} cos(\pi nx/L)e^{10t(\pi n/L)^2}=\sum B_n \pi n cos(\pi nx)e^{10t(\pi n)^2}"


"u_x(x,0)=\sum B_n \pi n cos(\pi nx)=x+1"


"B_n=20\int^1_0sin(\pi nx)dx=-\frac{20}{\pi n}cos(\pi nx)|^1_0=20/(\pi n)-20cos(\pi n)/(\pi n)"


"B_n=0" for even n

"B_n=40/(\pi n)" for odd n


"u_x(x,0)=40\sum cos(\pi nx)=x+1" , n=1,3,5,...


"u(x,t)=\sum\frac{40}{\pi n}sin(\pi nx)e^{10t(\pi n)^2}" , n=1,3,5,...